# Spectuner-D1: Spectral Line Fitting of Interstellar Molecules Using Deep Reinforcement Learning

### v1

@@ 3 Results
Figure 3 illustrates the relation between the relative fitting error and the number of forward modeling runs for different fitting methods. All the results are produced with $N_{\text{init}}=50$. We show the average metrics over all datasets, and the error bars represent the standard deviation estimated over the datasets. The training and testing datasets are summarized in Table 2 and described in Section 2.4. For local optimizers, using the neural network to generate initial guesses leads to significantly lower relative fitting loss compared to using random initial guesses. Some methods achieve results comparable to PSO with $N_{\text{trial}}=3$, while requiring significantly fewer forward modeling runs. Among the six local optimizers tested, the gradient-free Nelder-Mead algorithm achieves the lowest relative fitting loss of $0.01\pm 0.09$. The SLSQP algorithm performs best among gradient-based methods, achieving a relative fitting loss of $0.00\pm 0.09$. Additionally, the neural network can improve the performance of PSO. Using the neural network, PSO achieves better fitting results in a single run compared to PSO with $N_{\text{trial}}=3$ without neural network assistance. Furthermore, no significant decrease in the number of forward modeling runs is found when using the neural network. An explanation could be that the optimizers with random initial guesses tend to get stuck in local minima, and therefore stop earlier.
@@
We next evaluate the generalization ability of our neural network. Figure 5 demonstrates the mean relative loss for each dataset. As examples, we show only the results obtained using the neural network alone and the SLSQP algorithm. As described in Section 2.4, our testing data consists of two parts. First, the data cubes labeled ATOMS-20-band3-2as, IRAS16293 A-band3-1.0as, IRAS16293 B-band3-1.0as, and Orion KL-band6-4.9as are reserved for testing and were not used during training. Secondly, the neural network is trained using only species from Mol-1980, which includes molecules detected before 1980 along with their isotopologues, while we also carry out baseline fitting for species detected between 1980 and 2010, labeled as Mol-2010.
@@
4 Applying to pixel-level fitting
@@
We suggest that our neural network can be efficiently applied to pixel-level fitting. As an example, we apply our method to the central hot core region of G327-band6-1.9as. The region have $100\times 100$ pixels. We use 3 spectral windows, covering 215565 - 217273 MHz, 217572 - 219271 MHz and 229611 - 231319 MHz. This work performs fittings for each pixel independently, while our neural network based method can be applied to algorithms that take into account the correlation between the pixels. For instance, instead of fitting all pixels, 2025A&A...700A..84L introduced an active-learning framework to identify the most informative pixels for fitting, and approximated the parameter maps using the Gaussian-process model.
4.1 Results of methanol
@@
Figure 6 illustrates the pixel-level fitting results of CH3OH, $v=0$. The first row presents the results using random initial guesses, which are quite noisy. In contrast, as shown in the second row, the results using the neural network are more accurate. In the third and fourth rows, we compare our results with xclass (2017A&A...598A...7M). When producing the xclass results, we also provide the initial parameters given by our neural network. In the central hot core, our results are consistent with xclass.
@@
We then compare the results obtained using the $\chi^{2}$ and peak matching functions in Figure 7. While the predicted column densities are consistent within $\sim 0.5$ dex, the other parameters show discrepancies. The $v_{\text{offset}}$ distributions exhibit similarities, while the values have offsets. Compared to the $\chi^{2}$ results, the results using the peak matching loss tend to underestimate the velocity width. To demonstrate the problem, we also show the spectra of two pixels in Figure 7. For clarity, we only show the spectra over a representative frequency range. It can be found that the observed spectra (black lines) include features that our spectral line model cannot reproduce, e.g. the line wings. Our model assumes only a Gaussian line profile. The $\chi^{2}$ and peak matching loss functions fit the observed spectra in different ways. According to 2025ApJS..277...21Q, the peak matching algorithm only compares the top part of the peaks, while the $\chi^{2}$ function tends to fit the entire line shape. These may explain the discrepancies in the estimated parameters, in particular the velocity width and offset.
4.2 Results of more molecules
@@
In Figure 8, we present the fitting results of CH3OCHO, $v=0$, CH3OCH3, $v=0$, C2H5CN, $v=0$, and C2H3CN, $v=0$. These results were obtained by combining the neural network predictions with the SLSQP algorithm. The peak matching loss function is employed for the fittings. No meaningful results can be obtained for these COMs using $\chi^{2}$ fitting due to line blending. This issue was discussed in 2025ApJS..277...21Q.
@@
Our method achieves reasonable results for these COMs. Clear velocity gradients are visible in Figure 8. Both excitation temperatures and column densities increase toward the center. A detailed scientific explanation of these results is deferred to future studies, as this work focuses on demonstrating the fitting method.
@@
4.3 Performance
Whereas the training of our neural network requires dozens of GPUs as mentioned in Section 2.5, the inference only requires one consumer-grade GPU. The pixel-level fitting results presented above were performed using one RTX 2080 Ti card and 16 cores on a desktop. The fitting region has 10,000 pixels, with 8740 frequency channels. The execution time for pixel-by-pixel fitting of different species ranges from 4.9 minutes (29 ms/pixel) to 41.9 minutes (251 ms/pixel), as summarized in Table 5. For methanol, the runtime using the peak matching loss is long than using the $\chi^{2}$ loss. The peak matching loss is less smooth than the $\chi^{2}$ function and, on average, results in 2.4 times forward modeling runs during fitting. Its calculation is also slower than the $\chi^{2}$ function.
5 Discussion
This work trains a neural network to predict the five parameters of the one-dimensional LTE radiative transfer model. The model is described in Section 2.3.1. This model is suitable for COMs in star-forming regions with high H2 volume density (e.g 2020A&A...641A..54C; 2024A&A...686A.289H), where the LTE assumption holds. In addition, due to degeneracies among source size, excitation temperature, and column density, our method is unsuitable for those species with only a few emission lines in the spectrum, for instance CO and HCN.
@@ 5 Discussion
Furthermore, this study considers only molecular rotational spectra in the radio band, while molecules can also be detected via vibrational transitions in the infrared band using the James Webb Space Telescope (JWST) (e.g. 2023NatAs...7..431M; 2024A&A...690A.205C; 2024A&A...686A..71N). There are two main challenges in extending our methodology to infrared observations. First, our approach relies on the efficient generation of synthetic spectra for comparison with observations, while modeling infrared transitions is complex due to contributions from both gas and ice phases (2024A&A...690A.205C). Secondly, the baseline of infrared spectra is complex(e.g. 2023NatAs...7..431M; 2024A&A...690A.205C), whereas our peak-matching algorithm assumes a flat baseline. In such cases, local regression techniques or autoregressive moving average models (e.g. chatfield2019analysis; hyndman2021) could be used to correct baselines and extract spectral line features.
6 Summary
@@
This work presents a robust and efficient method for spectral line fitting of interstellar molecules, providing an excellent tool for analyzing data from large surveys, such as ATOMS (2020MNRAS.496.2790L), QUARKS (2024RAA....24b5009L), ALMA-IMF(2022A&A...662A...8M), and ALMAGAL (2025A&A...696A.149M).

### v2

@@ III Results
Figure 3 illustrates the relation between the relative fitting error and the number of forward modeling runs for different fitting methods. All the results are produced with $N_{\text{init}}=50$. We show the average metrics over all datasets, and the error bars represent the standard deviation estimated over the datasets. The training and testing datasets are summarized in Table 2 and described in Section II.4. For local optimizers, using the neural network to generate initial guesses leads to significantly lower relative fitting loss compared to using random initial guesses. Some methods achieve results comparable to PSO with $N_{\text{trial}}=3$, while requiring significantly fewer forward modeling runs. Among the six local optimizers tested, the gradient-free Nelder-Mead algorithm achieves the lowest relative fitting loss of $0.01\pm 0.09$. The SLSQP algorithm performs best among gradient-based methods, achieving a relative fitting loss of $0.00\pm 0.09$. Additionally, the neural network can improve the performance of PSO. Using the neural network, PSO achieves better fitting results in a single run compared to PSO with $N_{\text{trial}}=3$ without neural network assistance. Furthermore, no significant decrease in the number of forward modeling runs is found when using the neural network. An explanation could be that the optimizers with random initial guesses tend to get stuck in local minima, and therefore stop earlier.
@@
We next evaluate the generalization ability of our neural network. Figure 5 demonstrates the mean relative loss for each dataset. As examples, we show only the results obtained using the neural network alone and the SLSQP algorithm. As described in Section II.4, our testing data consists of two parts. First, the data cubes labeled ATOMS-20-band3-2as, IRAS16293 A-band3-1.0as, IRAS16293 B-band3-1.0as, and Orion KL-band6-4.9as are reserved for testing and were not used during training. Secondly, the neural network is trained using only species from Mol-1980, which includes molecules detected before 1980 along with their isotopologues, while we also carry out baseline fitting for species detected between 1980 and 2010, labeled as Mol-2010.
@@
IV Applying to pixel-level fitting
@@
We suggest that our neural network can be efficiently applied to pixel-level fitting. As an example, we apply our method to the central hot core region of G327-band6-1.9as. The region have $100\times 100$ pixels. We use 3 spectral windows, covering 215565 - 217273 MHz, 217572 - 219271 MHz and 229611 - 231319 MHz. This work performs fittings for each pixel independently, while our neural network based method can be applied to algorithms that take into account the correlation between the pixels. For instance, instead of fitting all pixels, Lin et al. (2025) introduced an active-learning framework to identify the most informative pixels for fitting, and approximated the parameter maps using the Gaussian-process model.
We perform pixel-by-pixel fitting for CH3OH, $v=0$, CH3OCHO, $v=0$, CH3OCH3, $v=0$, C2H5CN, $v=0$, and C2H3CN, $v=0$. These molecules are well-established tracers of hot core chemistry (e.g., Fontani et al., 2007; Li et al., 2024) and have been previously detected in the G327 hot core (Gibb et al., 2000).
@@
Figure 6 illustrates the pixel-level fitting results of CH3OH, $v=0$. The first row presents the results using random initial guesses, which are quite noisy. In contrast, as shown in the second row, the results using the neural network are more accurate. In the third and fourth rows, we compare our results with xclass (Möller et al., 2017). When producing the xclass results, we also provide the initial parameters given by our neural network. In the central hot core, our results are consistent with xclass.
@@
We then compare the results obtained using the $\chi^{2}$ and peak matching functions in Figure 7. While the predicted column densities are consistent within $\sim 0.5$ dex, the other parameters show discrepancies. The $v_{\text{offset}}$ distributions exhibit similarities, while the values have offsets. Compared to the $\chi^{2}$ results, the results using the peak matching loss tend to underestimate the velocity width. To demonstrate the problem, we also show the spectra of two pixels in Figure 7. For clarity, we only show the spectra over a representative frequency range. It can be found that the observed spectra (black lines) include features that our spectral line model cannot reproduce, e.g. the line wings. Our model assumes only a Gaussian line profile. The $\chi^{2}$ and peak matching loss functions fit the observed spectra in different ways. According to Qiu et al. (2025), the peak matching algorithm only compares the top part of the peaks, while the $\chi^{2}$ function tends to fit the entire line shape. These may explain the discrepancies in the estimated parameters, in particular the velocity width and offset.
@@
In Figure 8, we present the fitting results of CH3OCHO, $v=0$, CH3OCH3, $v=0$, C2H5CN, $v=0$, and C2H3CN, $v=0$. These results were obtained by combining the neural network predictions with the SLSQP algorithm. The peak matching loss function is employed for the fittings. No meaningful results can be obtained for these COMs using $\chi^{2}$ fitting due to line blending. This issue was discussed in Qiu et al. (2025).
@@
Our method achieves reasonable results for these COMs. Clear velocity gradients are visible in Figure 8. Both excitation temperatures and column densities increase toward the center. These trends imply rotational motion and trace the physical structure of the G327 hot core. A detailed scientific explanation of these results is deferred to future studies, as this work focuses on demonstrating the fitting method.
@@
IV.1 Performance
Whereas the training of our neural network requires dozens of GPUs as mentioned in Section II.5, the inference only requires one consumer-grade GPU. The pixel-level fitting results presented above were performed using one RTX 2080 Ti card and 16 cores on a desktop. The fitting region has 10,000 pixels, with 8740 frequency channels. The execution time for pixel-by-pixel fitting of different species ranges from 4.9 minutes (29 ms/pixel) to 41.9 minutes (251 ms/pixel), as summarized in Table 5. For methanol, the runtime using the peak matching loss is long than using the $\chi^{2}$ loss. The peak matching loss is less smooth than the $\chi^{2}$ function and, on average, results in 2.4 times forward modeling runs during fitting. Its calculation is also slower than the $\chi^{2}$ function.
V Discussion
This work trains a neural network to predict the five parameters of the one-dimensional LTE radiative transfer model. The model is described in Section II.3.1. This model is suitable for COMs in star-forming regions with high H2 volume density (e.g Coletta et al., 2020; Hsieh et al., 2024), where the LTE assumption holds. In addition, due to degeneracies among source size, excitation temperature, and column density, our method is unsuitable for those species with only a few emission lines in the spectrum, for instance CO and HCN.
@@ V Discussion
Furthermore, this study considers only molecular rotational spectra in the radio band, while molecules can also be detected via vibrational transitions in the infrared band using the James Webb Space Telescope (JWST) (e.g. McClure et al., 2023; Chen et al., 2024; Nazari et al., 2024b). There are two main challenges in extending our methodology to infrared observations. First, our approach relies on the efficient generation of synthetic spectra for comparison with observations, while modeling infrared transitions is complex due to contributions from both gas and ice phases (Chen et al., 2024). Secondly, the baseline of infrared spectra is complex(e.g. McClure et al., 2023; Chen et al., 2024), whereas our peak-matching algorithm assumes a flat baseline. In such cases, local regression techniques or autoregressive moving average models (e.g. Chatfield and Xing, 2019; Hyndman and Athanasopoulos, 2021) could be used to correct baselines and extract spectral line features.
VI Summary
@@
This work presents a robust and efficient method for spectral line fitting of interstellar molecules, providing an excellent tool for analyzing data from large surveys, such as ATOMS (Liu et al., 2020), QUARKS (Liu et al., 2024), ALMA-IMF(Motte et al., 2022), and ALMAGAL (Molinari et al., 2025).